Given f(x) = –2x + 4, find f(a – 2).
Answer by funmath(2933)

Given f(x) = –2x + 4, find f(a – 2).
Replace x with (a-2) in f(x):
f(a-2)=-2(a-2)+4
f(a-2)=-2(a)-2(-2)+4
f(a-2)=-2a+4+4
f(a-2)=-2a+8
:
Happy Calculating!!!